OpenStudy (anonymous):

I need 5 Calculus ii problems that have to not consist of any fractions, decimals, or odds.

OpenStudy (p0sitr0n):

you need exercises?

OpenStudy (anonymous):

No just the problems. Any kind of problem only if it does not have decimals fractions or odds.

OpenStudy (p0sitr0n):

evaluate\[\int\limits_{-\infty}^{\infty}\frac{ 1 }{ 1+x^{2}} dx\]

OpenStudy (acxbox22):

it has a fraction

OpenStudy (anonymous):

I need four more, I will alter that problem.

OpenStudy (p0sitr0n):

the answer is not a decimal or an odd.

OpenStudy (anonymous):

The equation has fraction and odd

OpenStudy (p0sitr0n):

\[\int\limits_{0}^{1}x^{3}\sqrt{1-x ^{2}}dx\]

OpenStudy (p0sitr0n):

you can find the integral of secant x (sec(x))

OpenStudy (anonymous):

Do you have 3 more?

OpenStudy (p0sitr0n):

for a challenge: \[\int\limits_{?}^{?} \frac{ 1 }{ 1+x ^{4} }dx\]

OpenStudy (p0sitr0n):

volumes: using either methods, compute the voly

OpenStudy (p0sitr0n):

volume under f(x) =e^x from 2 to - infinity

OpenStudy (p0sitr0n):

also for fun you could prove the fundamental theorem of calculus, part I or II, doesnt involve any decimals or fractions lol

OpenStudy (anonymous):

Thank you P0sitr0n :)